# Bounded Hyperbolic Tangent: A Stable and Efficient Alternative to Pre-Layer Normalization in Large Language Models

### v1

@@ Abstract
Pre-Layer Normalization (Pre-LN) is the de facto choice for large language models (LLMs) and is crucial for stable pretraining and effective transfer learning. However, Pre-LN is inefficient due to repeated statistical calculations and suffers from the curse of depth. As layers grow, the magnitude and variance of the hidden state escalate, destabilizing training. Efficiency-oriented normalization-free methods such as Dynamic Tanh (DyT) improve speed but remain fragile at depth. To jointly address stability and efficiency, we propose Bounded Hyperbolic Tanh (BHyT), a drop-in replacement for Pre-LN. BHyT couples a tanh nonlinearity with explicit, data-driven input bounding to keep activations within a non-saturating range. It prevents depth-wise growth in activation magnitude and variance and comes with a theoretical stability guarantee. For efficiency, BHyT computes exact statistics once per block and replaces a second normalization with a lightweight variance approximation, enhancing efficiency. Empirically, BHyT demonstrates improved stability and efficiency during pretraining, achieving an average of 15.8% faster training and an average of 4.2% higher token generation throughput compared to RMSNorm., while matching or surpassing its inference performance and robustness across language understanding and reasoning benchmarks. Our code is available at: https://anonymous.4open.science/r/BHyT
@@ 1 Introduction
The remarkable progress of large language models (LLMs) is tightly coupled to the Transformer architecture, where normalization layers play a central role vaswani2017attention. In particular, Pre‑Layer Normalization (Pre‑LN), which applies normalization before each sublayer, has become the de facto standard xiong2020layer; radford2019language; brown2020language; touvron2023llama. This design stabilizes optimization in very deep networks, enabling training at scale and strong transfer performance across diverse tasks. By regulating the scale and distribution of activations, Pre‑LN provides a reliable training regime that helps unlock the modern era of large‑scale LLMs ba2016layer; xiong2020layer.
As models push to greater depths, however, these foundation models expose a critical weakness. Recent analyses identify a curse of depth, in which the interaction between residual connections and Pre‑LN drives rapid, often near‑exponential, growth in hidden‑state magnitude and variance with layer depth (sun2025curse). This escalation is more important than simple divergence: it drives the block Jacobian toward the identity, so deep layers learn little beyond an expensive identity mapping sun2025curse. The core challenge is thus not merely preventing crashes; it is preserving informative transformations and efficient signal propagation so that every layer contributes meaningfully to learning.
The community has explored remedies along a stability–efficiency axis. On the stability end, stronger normalization seeks to suppress depth‑wise drift. For instance, Peri‑LN applies normalization both before and after each sublayer, curbing variance growth and maintaining gradient health even at extreme depths kim2025periln. The trade‑off is substantial overhead: doubled normalization increases reduction operations and memory traffic, inflating latency. On the efficiency end, RMSNorm removes mean centering to reduce computational costs and has been widely adopted in modern LLMs zhang2019root. Yet it only partially controls activation statistics and does not fundamentally resolve variance accumulation, leaving depth scaling fragile.
@@
A more radical line of work questions the need for normalization at all. Dynamic Tanh (DyT) replaces normalization with a simple element‑wise nonlinearity and a learnable scale zhu2025transformers. By eliminating per‑token statistics, DyT accelerates both training and inference. However, this efficiency comes with a major risk: activation saturation. When inputs grow large, precisely the hazard under depth, the nonlinearity saturates, gradients vanish, and optimization becomes brittle. Relying on a learned global scale to self‑regulate these dynamics is often insufficient in the face of strong depth‑induced amplification.
This landscape reveals a clear gap: methods that guarantee stability typically sacrifice efficiency, while efficient methods fail to robustly address the curse of depth. What is needed is a mechanism that retains the simplicity and throughput of normalization‑free designs while providing principled control over depth‑wise growth.
We introduce Bounded Hyperbolic Tanh (BHyT) to close this gap. As shown in Figure 1, BHyT is a drop‑in replacement for Pre‑LN that couples a tanh nonlinearity with explicit, data‑driven input bounding. By keeping activations within a non-saturating regime with high probability, BHyT prevents depth-wise growth in activation magnitude and variance, preserving informative transformations as depth increases. Beyond this forward‑path protection, BHyT comes with a theoretical stability guarantee: its gradient scale is provably upper‑bounded by that of RMSNorm, ensuring at least comparable stability even in very deep stacks. To minimize overhead, BHyT computes exact statistics once per block and replaces a second normalization with a lightweight variance approximation that tracks attention‑induced variation, reducing reductions and memory movement. The result is a practical module that maintains stability without reintroducing the throughput penalties of heavy normalization.
@@ 2.1 Limitations of Pre-Layer Normalization
Layer normalization has been a key component in Transformer architectures (ba2016layer), and its Pre-LN variant has become the default in most LLMs (xiong2020layer). The Pre-LN design, which normalizes hidden states before attention and feed-forward layers, improves stability when training begins. Still, this design is not without drawbacks. From a stability perspective, recent studies have shown that Pre-LN suffers from massive activations as depth increases (sun2025curse; kim2025periln). In this regime, the mean and variance of hidden states grow rapidly across layers, destabilizing optimization. From an efficiency perspective, normalization requires computing statistics such as the mean and variance of activations. These reductions introduce latency and memory overhead, which accumulate as depth increases. As a result, Pre-LN can both destabilize very deep models and become a bottleneck for training throughput and scalability.
@@ 2.2 Normalization-Based Enhancements
The main bottleneck of normalization-based designs lies in the repeated computation of statistics at every block. A first attempt to reduce this cost was RMSNorm, which discards the mean and normalizes only by the root mean square of activations (zhang2019root). This removes part of the overhead while keeping activations within a reasonable range. However, variance computation is still required for every layer, and the instability of Pre-LN remains unresolved.
To address this instability, LayerNorm Scaling (LNS) (sun2025curse) introduces a layer-index scaling that constrains variance growth across depth. By rescaling activations according to layer index, it alleviates the massive activation problem and stabilizes optimization in very deep networks. Although LNS improves stability, its reliance on per-block statistics keeps the cost high in large-scale training. Peri-LN reinforces stability further by applying normalization both before and after each sublayer (kim2025periln). This suppresses variance spikes and improves convergence and downstream accuracy; however, doubling normalization operations adds substantial overhead, making it less suitable for large-scale pretraining.
@@ 2.3 Normalization-Free Alternatives
To avoid the overhead of statistical computations altogether, recent work has explored removing normalization. DyT follows this direction by replacing RMSNorm with a bounded activation function, $\tanh(\alpha_{\text{DyT}}x)$, where $\alpha_{\text{DyT}}$ is a learnable scalar (zhu2025transformers). This eliminates the need for mean or variance estimation, improving both training and inference speeds. The $\tanh$ function ensures that activations remain bounded, and the learnable scalar $\alpha_{\text{DyT}}$ rescales inputs so that the effective range of $\tanh$ is better utilized.
@@ 3 Methodology
An ideal normalization scheme should allow for controlled, sub-exponential or linear growth of variance (sun2025curse; kim2025periln), sufficient to preserve signal propagation without leading to explosive instability. To this end, we introduce BHyT, a novel normalization layer designed to achieve both stability and speed. BHyT builds upon the efficiency of methods like DyT, which use S-shaped functions (e.g., $\tanh$) to bypass explicit statistical computation. However, DyT’s reliance on a single learnable scalar to control the input to $\tanh$ offers no explicit guarantee against saturation, which can lead to vanishing gradients in very deep models. BHyT addresses this limitation by explicitly bounding the input to the non-linear function, ensuring both numerical stability and a well-behaved gradient flow.
@@ 3.1 Ensuring Stability via Input Bounding
Specifically, we utilize Chebyshev’s inequality, which guarantees that for any random variable $X$ with a finite mean $\mu$ and variance $s^{2}$, the probability of deviating from the mean by less than $\kappa$ standard deviations is at least $1-1/\kappa^{2}$ (i.e., $P(|X-\mu|<\kappa s)\geq 1-\kappa^{-2}$).
@@ 3.2.2 BHyT before the MLP Layer
Let the sequence length $T$ be sufficiently large so that the attention weights can be treated as approximately uniform (ledoux2001concentration). Let the token-level activation $x\in\mathbb{R}^{d}$ and the projection matrices $W_{V}$ and $W_{O}$ follow independent zero-mean normal distributions. By these assumptions, BHyT keeps the argument of the $\tanh(\cdot)$ function in its near-linear regime, such that scaling by $\lambda/\kappa s_{x}$ directly yields a variance of approximately $\lambda^{2}/\kappa^{2}$.
@@ 3.3 Depth-wise Variance Propagation and Stability of BHyT
For a network of depth $L$, if the BHyT hyperparameters satisfy $\frac{\lambda}{\kappa}<\frac{1}{\sqrt{L}}$, then the output variance of BHyT is strictly smaller than that of LNS (sun2025curse) for every layer $\ell$ up to depth $L$ (i.e., for all $1\leq\ell\leq L$):
@@ 4 Experiments & results
In this section, we present a comprehensive set of experiments designed to evaluate the effectiveness of our proposed method, BHyT. Our evaluation is structured around three primary objectives: 1) Stability, assessing BHyT’s ability to mitigate the “curse of depth” (sun2025curse) by analyzing activation growth across layers; 2) Efficiency, comparing its computational throughput against existing normalization methods; and 3) Performance, measuring its effectiveness in both pretraining and supervised fine-tuning (SFT) contexts to ensure it achieves competitive results.
@@ 4.1 Experimental Setup
We conduct experiments with two model scales from the Llama-3.2 family: a 1B-parameter model with 16 layers and a 3B-parameter model with 28 layers. Rather than using pretrained checkpoints, we instantiate only the architectures and train all models from scratch. This setup allows us to explicitly examine the effect of increasing model depth on stability and efficiency and to highlight the scalability benefits of our proposed BHyT method.
@@
Pretraining is performed on the C4 corpus, a large-scale English text dataset widely used in language model pretraining (raffel2020exploring). For supervised fine-tuning (SFT), we employ an instruction-tuning dataset: Lima1k (zhou2023lima), which contains carefully curated, high-quality instructions that provide broader coverage with a larger number of instruction–response pairs. This combination enables us to evaluate both high-quality, low-resource, and large-scale instruction-tuning scenarios.
Following pretraining and SFT, we assess model performance on a diverse suite of seven language understanding and reasoning benchmarks: ARC-e (clark2018think), PIQA (bisk2020piqa), HellaSwag (zellers2019hellaswag), OpenBookQA (OpenBookQA2018), Winogrande (sakaguchi2021winogrande), MMLU (hendrycks2020measuring), and BoolQ (clark-etal-2019-boolq). These benchmarks collectively evaluate both factual knowledge and reasoning abilities. We report task-specific accuracy and macro-average accuracy across tasks as the primary evaluation metric.
@@
All pretraining and SFT experiments are implemented using the Llama Factory framework (zheng2024llamafactory), which provides efficient fine-tuning utilities and ensures consistent implementation across methods. For downstream evaluation, we use the lm-evaluation-harness toolkit (eval-harness), adopting its standardized task configurations to ensure comparability with prior work. Accuracy scores are computed directly by the harness without modification, ensuring the reproducibility of reported results.
@@ 4.2 Pretraining LLMs
We compare BHyT against several baselines: RMSNorm (zhang2019root), the default normalization used in the Llama family; LNS (sun2025curse) and Peri-LN (kim2025periln), both designed to enhance training stability; and DyT (zhu2025transformers), which replaces the conventional Pre-LN with a $\tanh(\cdot)$-based function having learnable scalar parameter $\alpha_{\text{DyT}}$. We conducted pretraining on the C4 dataset. The 1B model was optimized with a sequence length of 1024, a batch size of 32, and for 50,000 steps, amounting to $\sim$1.64B tokens. The 3B model used the same configuration but was trained for 60,000 steps, resulting in $\sim$1.97B tokens. To further evaluate each method under a more realistic large-scale setting, we additionally extend the pretraining budget to 20B tokens and conduct experiments on both the 1B and 3B models.
@@
Table 1: Pretraining (PT)-only evaluation for Llama-1B and Llama-3B; downstream benchmarks in a 5-shot setting, averaged over five training seeds. The results demonstrate that BHyT achieves performance comparable to or superior to strong baselines while maintaining high computational speed. PPL denotes the perplexity score.
Llama-1B
Method
PT Train
Loss
PT Eval
Loss
PT Eval
PPL
Arc-e PIQA Hellaswag OpenBookQA Winogrande MMLU BoolQ Avg.
[…]
@@
RMSNorm 3.203 3.180 24.040 $31.01_{\pm{0.21}}$ $\textbf{66.57}_{\pm{0.41}}$ $\textbf{41.52}_{\pm{0.25}}$ $\textbf{33.92}_{\pm{0.33}}$ $50.50_{\pm{0.78}}$ $25.67_{\pm{0.02}}$ $37.94_{\pm{0.12}}$ $41.02_{\pm{0.18}}$
@@
LNS 3.160 3.139 23.091 $\textbf{31.88}_{\pm{0.42}}$ $64.68_{\pm{0.46}}$ $36.25_{\pm{0.10}}$ $\textit{32.45}_{\pm{0.53}}$ $\textit{51.18}_{\pm{0.75}}$ $\textbf{26.87}_{\pm{0.01}}$ $37.83_{\pm{0.08}}$ $40.16_{\pm{0.11}}$
@@
BHyT 3.133 3.107 22.346 $\textit{31.84}_{\pm{0.15}}$ $\textit{65.08}_{\pm{0.39}}$ $\textit{36.48}_{\pm{0.04}}$ $31.76_{\pm{0.33}}$ $\textbf{51.62}_{\pm{0.66}}$ $25.70_{\pm{0.01}}$ $\textbf{60.84}_{\pm{0.29}}$ $\textbf{43.44}_{\pm{0.16}}$
@@
4.3 Instruction Fine-tuning and Benchmark Evaluation
Table 2: Supervised fine-tuning (SFT) results across five training seeds for Llama-1B and Llama-3B. Reported values include SFT losses and 5-shot downstream benchmark accuracies; BHyT attains lower SFT loss and competitive performance, indicating effective transfer of its pretraining stability to instruction-tuned settings.
Llama-1B
Method
PT Eval
Loss
SFT Train
Loss
SFT Eval
Loss
Arc-e PIQA Hellaswag OpenBookQA Winogrande MMLU BoolQ Avg.
@@ 4.3 Instruction Fine-tuning and Benchmark Evaluation
RMSNorm 3.272 $\textit{2.646}_{\pm{0.011}}$ $3.217_{\pm{0.130}}$ $\textbf{37.67}_{\pm{0.23}}$ $\textit{66.96}_{\pm{0.29}}$ $31.70_{\pm{0.12}}$ $\textit{31.12}_{\pm{0.18}}$ $\textbf{51.22}_{\pm{1.60}}$ $25.99_{\pm{0.04}}$ $53.31_{\pm{2.02}}$ $39.83_{\pm{0.21}}$
Peri-LN 3.279 $\textbf{2.614}_{\pm{0.011}}$ $3.178_{\pm{0.132}}$ $\textit{36.80}_{\pm{0.27}}$ $\textbf{67.05}_{\pm{0.12}}$ $\textbf{37.13}_{\pm{0.09}}$ $\textbf{31.44}_{\pm{0.61}}$ $49.06_{\pm{0.40}}$ $\textbf{26.36}_{\pm{0.02}}$ $52.75_{\pm{1.54}}$ $\textit{42.45}_{\pm{0.29}}$
LNS 3.271 $2.652_{\pm{0.011}}$ $\textit{3.157}_{\pm{0.132}}$ $34.49_{\pm{0.17}}$ $65.59_{\pm{0.16}}$ $28.96_{\pm{0.54}}$ $28.96_{\pm{0.54}}$ $50.61_{\pm{0.55}}$ $\textit{26.33}_{\pm{0.02}}$ $52.69_{\pm{0.66}}$ $41.78_{\pm{0.20}}$
DyT 3.855 $3.361_{\pm{0.010}}$ $3.971_{\pm{0.132}}$ $29.78_{\pm{0.40}}$ $58.42_{\pm{0.53}}$ $25.95_{\pm{0.17}}$ $28.28_{\pm{0.90}}$ $49.53_{\pm{0.79}}$ $25.82_{\pm{0.13}}$ $\textit{56.64}_{\pm{1.02}}$ $38.87_{\pm{0.17}}$
BHyT 3.254 $2.693_{\pm{0.012}}$ $\textbf{3.130}_{\pm{0.133}}$ $34.61_{\pm{0.20}}$ $66.47_{\pm{0.27}}$ $\textit{36.95}_{\pm{0.12}}$ $29.68_{\pm{0.46}}$ $\textit{51.14}_{\pm{1.02}}$ $26.06_{\pm{0.05}}$ $\textbf{58.21}_{\pm{0.81}}$ $\textbf{42.88}_{\pm{0.19}}$
We apply LoRA-based supervised fine-tuning to assess whether the stability achieved during pretraining carries over to instruction-tuned performance. Table 2 summarizes the SFT losses and downstream benchmark accuracies for the Llama-1B and Llama-3B. Across both scales, BHyT maintains the stable optimization observed during pretraining and transfers this advantage effectively to the SFT stage. Notably, BHyT attains lower SFT loss and competitive accuracy on a broad set of benchmarks, demonstrating that its depth-wise stability and controlled activation dynamics continue to support reliable adaptation beyond pretraining.

### v2

@@ Abstract
Pre-Layer Normalization (Pre-LN) is the de facto choice for large language models (LLMs) and is crucial for stable pretraining and effective transfer learning. However, Pre-LN is inefficient due to repeated statistical calculations and suffers from the curse of depth. As layers grow, the magnitude and variance of the hidden state escalate, destabilizing training. Efficiency-oriented normalization-free methods such as Dynamic Tanh (DyT) improve speed but remain fragile at depth. To jointly address stability and efficiency, we propose Bounded Hyperbolic Tanh (BHyT), a drop-in replacement for Pre-LN. BHyT couples a tanh nonlinearity with explicit, data-driven input bounding to keep activations within a non-saturating range. It prevents depth-wise growth in activation magnitude and variance and comes with a theoretical stability guarantee. For efficiency, BHyT computes exact statistics once per block and replaces a second normalization with a lightweight variance approximation, enhancing efficiency. Empirically, BHyT demonstrates improved stability and efficiency during pretraining, achieving an average of 15.8% faster training and an average of 4.2% higher token generation throughput compared to RMSNorm, while matching or surpassing its inference performance and robustness across language understanding and reasoning benchmarks. Our code is available at: https://anonymous.4open.science/r/BHyT
@@ 1 Introduction
The remarkable progress of large language models (LLMs) is tightly coupled to the Transformer architecture, where normalization layers play a central role Vaswani et al. (2017). In particular, Pre‑Layer Normalization (Pre‑LN), which applies normalization before each sublayer, has become the de facto standard Xiong et al. (2020); Radford et al. (2019); Brown et al. (2020); Touvron et al. (2023). This design stabilizes optimization in very deep networks, enabling training at scale and strong transfer performance across diverse tasks. By regulating the scale and distribution of activations, Pre‑LN provides a reliable training regime that helps unlock the modern era of large‑scale LLMs Ba et al. (2016); Xiong et al. (2020).
As models push to greater depths, however, these foundation models expose a critical weakness. Recent analyses identify a curse of depth, in which the interaction between residual connections and Pre‑LN drives rapid, often near‑exponential, growth in hidden‑state magnitude and variance with layer depth (Sun et al., 2025). This escalation is more important than simple divergence: it drives the block Jacobian toward the identity, so deep layers learn little beyond an expensive identity mapping Sun et al. (2025). The core challenge is thus not merely preventing crashes; it is preserving informative transformations and efficient signal propagation so that every layer contributes meaningfully to learning.
The community has explored remedies along a stability–efficiency axis. On the stability end, stronger normalization seeks to suppress depth‑wise drift. For instance, Peri‑LN applies normalization both before and after each sublayer, curbing variance growth and maintaining gradient health even at extreme depths Kim et al. (2025). The trade‑off is substantial overhead: doubled normalization increases reduction operations and memory traffic, inflating latency. On the efficiency end, RMSNorm removes mean centering to reduce computational costs and has been widely adopted in modern LLMs Zhang and Sennrich (2019). Yet it only partially controls activation statistics and does not fundamentally resolve variance accumulation, leaving depth scaling fragile.
@@
A more radical line of work questions the need for normalization at all. Dynamic Tanh (DyT) replaces normalization with a simple element‑wise nonlinearity and a learnable scale Zhu et al. (2025). By eliminating per‑token statistics, DyT accelerates both training and inference. However, this efficiency comes with a major risk: activation saturation. When inputs grow large, precisely the hazard under depth, the nonlinearity saturates, gradients vanish, and optimization becomes brittle. Relying on a learned global scale to self-regulate these dynamics is often insufficient in the face of strong depth‑induced amplification.
This landscape reveals a clear gap: methods that guarantee stability typically sacrifice efficiency, while efficient methods fail to robustly address the curse of depth (Sun et al., 2025). What is needed is a mechanism that retains the simplicity and throughput of normalization‑free designs while providing principled control over depth‑wise growth.
We introduce Bounded Hyperbolic Tanh (BHyT) to close this gap. As shown in Figure 1, BHyT is a drop‑in replacement for Pre‑LN that couples a tanh nonlinearity with explicit, data‑driven input bounding. By keeping activations within a non-saturating regime with high probability, BHyT prevents depth-wise growth in activation magnitude and variance, preserving informative transformations as depth increases. Beyond this forward‑path protection, BHyT provides a theoretical stability guarantee regarding variance propagation. Specifically, for a network of depth $L$, the output variance of BHyT is provably smaller than that of LayerNorm Scaling (LNS) (Sun et al., 2025) under specific hyperparameter conditions. This ensures that BHyT more effectively suppresses depth‑wise variance growth compared to existing stable normalization methods, thereby mitigating the curse of depth (Sun et al., 2025) even in very deep architectures. To minimize overhead, BHyT computes exact statistics once per block and replaces a second normalization with a lightweight variance approximation that tracks attention‑induced variation, reducing reductions and memory movement. The result is a practical module that maintains stability without reintroducing the throughput penalties of heavy normalization.
@@ 2.1 Limitations of Pre-Layer Normalization
Layer normalization has been a key component in Transformer architectures (Ba et al., 2016), and its Pre-LN variant has become the default in most LLMs (Xiong et al., 2020). The Pre-LN design, which normalizes hidden states before attention and feed-forward layers, improves stability when training begins. Still, this design is not without drawbacks. From a stability perspective, recent studies have shown that Pre-LN suffers from massive activations as depth increases (Sun et al., 2025; Kim et al., 2025). In this regime, the mean and variance of hidden states grow rapidly across layers, destabilizing optimization. From an efficiency perspective, normalization requires computing statistics such as the mean and variance of activations. These reductions introduce latency and memory overhead, which accumulate as depth increases. As a result, Pre-LN can both destabilize very deep models and become a bottleneck for training throughput and scalability.
@@ 2.2 Normalization-Based Enhancements
The main bottleneck of normalization-based designs lies in the repeated computation of statistics at every block. A first attempt to reduce this cost was RMSNorm, which discards the mean and normalizes only by the root mean square of activations (Zhang and Sennrich, 2019). This removes part of the overhead while keeping activations within a reasonable range. However, variance computation is still required for every layer, and the instability of Pre-LN remains unresolved.
To address this instability, LayerNorm Scaling (LNS) (Sun et al., 2025) introduces a layer-index scaling that constrains variance growth across depth. By rescaling activations according to layer-index, it alleviates the massive activation problem and stabilizes optimization in very deep networks. Although LNS improves stability, its reliance on per-block statistics keeps the cost high in large-scale training. Peri-LN reinforces stability further by applying normalization both before and after each sublayer (Kim et al., 2025). This suppresses variance spikes and improves convergence and downstream accuracy; however, doubling normalization operations adds substantial overhead, making it less suitable for large-scale pretraining.
@@ 2.3 Normalization-Free Alternatives
To avoid the overhead of statistical computations altogether, recent work has explored removing normalization. DyT follows this direction by replacing RMSNorm with a bounded activation function, $\tanh(\alpha_{\text{DyT}}x)$, where $\alpha_{\text{DyT}}$ is a learnable scalar (Zhu et al., 2025). This eliminates the need for mean or variance estimation, improving both training and inference speeds. The $\tanh$ function ensures that activations remain bounded, and the learnable scalar $\alpha_{\text{DyT}}$ rescales inputs so that the effective range of $\tanh$ is better utilized.
@@ 3 Methodology
An ideal normalization scheme should allow for controlled, sub-exponential or linear growth of variance (Sun et al., 2025; Kim et al., 2025), sufficient to preserve signal propagation without leading to explosive instability. To this end, we introduce BHyT, a novel normalization layer designed to achieve both stability and speed. BHyT builds upon the efficiency of methods like DyT, which use S-shaped functions (e.g., $\tanh$) to bypass explicit statistical computation. However, DyT’s reliance on a single learnable scalar to control the input to $\tanh$ offers no explicit guarantee against saturation, which can lead to vanishing gradients in very deep models. BHyT addresses this limitation by explicitly bounding the input to the nonlinear function, ensuring both numerical stability and a well-behaved gradient flow.
@@ 3.1 Ensuring Stability via Input Bounding
Specifically, we utilize Chebyshev’s inequality, which guarantees that for any random variable $X$ with a finite mean $\mu$ and variance $s^{2}$, the probability of deviating from the mean by less than $\kappa$ standard deviations is at least $1-1/\kappa^{2}$ (i.e., $P(|X-\mu|\leq\kappa s)\geq 1-\kappa^{-2}$).
@@ 3.2.2 BHyT before the MLP Layer
Let the sequence length $T$ be sufficiently large so that the attention weights can be treated as approximately uniform (Ledoux, 2001). Let the token-level activation $x\in\mathbb{R}^{d}$ and the projection matrices $W_{V}$ and $W_{O}$ follow independent zero-mean normal distributions. By these assumptions, BHyT keeps the argument of the $\tanh(\cdot)$ function in its near-linear regime, such that scaling by $\lambda/\kappa s_{x}$ directly yields a variance of approximately $\lambda^{2}/\kappa^{2}$.
@@ 3.3 Depth-wise Variance Propagation and Stability of BHyT
For a network of depth $L$, if the BHyT hyperparameters satisfy $\frac{\lambda}{\kappa}<\frac{1}{\sqrt{L}}$, then the output variance of BHyT is strictly smaller than that of LNS (Sun et al., 2025) for every layer $\ell$ up to depth $L$ (i.e., for all $1\leq\ell\leq L$):
@@ 4 Experiments & Results
In this section, we present a comprehensive set of experiments designed to evaluate the effectiveness of our proposed method, BHyT. Our evaluation is structured around three primary objectives: 1) Stability, assessing BHyT’s ability to mitigate the “curse of depth” (Sun et al., 2025) by analyzing activation growth across layers; 2) Efficiency, comparing its computational throughput against existing normalization methods; and 3) Performance, measuring its effectiveness in both pretraining and supervised fine-tuning (SFT) contexts to ensure it achieves competitive results.
@@ 4.1 Experimental Setup
We conduct experiments with two model scales from the Llama-3.2 family: a 1B-parameter model with 16 layers and a 3B-parameter model with 28 layers. We instantiate only the architectures and train all models from scratch. This setup allows us to explicitly examine the effect of increasing model depth on stability and efficiency and to highlight the scalability benefits of our proposed BHyT method.
@@
Pretraining is performed on the C4 corpus, a large-scale English text dataset widely used in language model pretraining (Raffel et al., 2020). For supervised fine-tuning (SFT), we employ an instruction-tuning dataset: Lima1k (Zhou et al., 2023), which contains carefully curated, high-quality instructions that provide broader coverage with a larger number of instruction–response pairs. This combination enables us to evaluate both high-quality, low-resource, and large-scale instruction-tuning scenarios.
Following pretraining and SFT, we assess model performance on a diverse suite of seven language understanding and reasoning benchmarks: ARC-e (Clark et al., 2018), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), Winogrande (Sakaguchi et al., 2021), MMLU (Hendrycks et al., 2020), and BoolQ (Clark et al., 2019). These benchmarks collectively evaluate both factual knowledge and reasoning abilities. We report task-specific accuracy and macro-average accuracy across tasks as the primary evaluation metric.
@@
All pretraining and SFT experiments are implemented using the Llama Factory framework (Zheng et al., 2024), which provides efficient fine-tuning utilities and ensures consistent implementation across methods. For downstream evaluation, we use the lm-evaluation-harness toolkit (Gao et al., 2024), adopting its standardized task configurations to ensure comparability with prior work. Accuracy scores are computed directly by the harness without modification, ensuring the reproducibility of reported results.
@@ 4.2 Pretraining LLMs
We compare BHyT against several baselines: RMSNorm (Zhang and Sennrich, 2019), the default normalization used in the Llama family; LNS (Sun et al., 2025) and Peri-LN (Kim et al., 2025), both designed to enhance training stability; and DyT (Zhu et al., 2025), which replaces the conventional Pre-LN with a $\tanh(\cdot)$-based function having learnable scalar parameter $\alpha_{\text{DyT}}$. We conducted pretraining on the C4 dataset. The 1B model was optimized with a sequence length of 1024, a batch size of 32, and for 50,000 steps, amounting to $\sim$1.64B tokens. The 3B model used the same configuration but was trained for 60,000 steps, resulting in $\sim$1.97B tokens. To further evaluate each method under a more realistic large-scale setting, we additionally extend the pretraining budget to 20B tokens and conduct experiments on both the 1B and 3B models.
@@
Table 1: Pretraining (PT)-only evaluation for Llama-3B; downstream benchmarks in a 5-shot setting, averaged over five training seeds. The results demonstrate that BHyT achieves performance comparable to or superior to strong baselines while maintaining high computational speed. PPL denotes the perplexity score.
@@
RMSNorm 3.203 3.180 24.040 $31.01_{\pm{0.21}}$ $\textbf{66.57}_{\pm{0.41}}$ $\textbf{41.52}_{\pm{0.25}}$ $\textbf{33.92}_{\pm{0.33}}$ $50.50_{\pm{0.78}}$ $25.67_{\pm{0.02}}$ $37.94_{\pm{0.12}}$ $40.89_{\pm{0.18}}$
@@
LNS 3.160 3.139 23.091 $\textbf{31.88}_{\pm{0.42}}$ $64.68_{\pm{0.46}}$ $36.25_{\pm{0.10}}$ $\textit{32.45}_{\pm{0.53}}$ $\textit{51.18}_{\pm{0.75}}$ $\textbf{26.87}_{\pm{0.01}}$ $37.83_{\pm{0.08}}$ $41.16_{\pm{0.11}}$
@@
BHyT 3.133 3.107 22.346 $\textit{31.84}_{\pm{0.15}}$ $\textit{65.08}_{\pm{0.39}}$ $\textit{36.48}_{\pm{0.04}}$ $31.76_{\pm{0.33}}$ $\textbf{51.62}_{\pm{0.66}}$ $25.70_{\pm{0.01}}$ $\textbf{60.84}_{\pm{0.29}}$ $\textbf{43.33}_{\pm{0.16}}$
@@
Table 2: Supervised fine-tuning (SFT) results across five training seeds for Llama-3B. Reported values include SFT losses and 5-shot downstream benchmark accuracies; BHyT attains lower SFT loss and competitive performance, indicating effective transfer of its pretraining stability to instruction-tuned settings.
@@
RMSNorm 3.272 $\textit{2.646}_{\pm{0.011}}$ $3.217_{\pm{0.130}}$ $\textbf{37.67}_{\pm{0.23}}$ $\textit{66.96}_{\pm{0.29}}$ $31.70_{\pm{0.12}}$ $\textit{31.12}_{\pm{0.18}}$ $\textbf{51.22}_{\pm{1.60}}$ $22.95_{\pm{0.04}}$ $53.31_{\pm{2.02}}$ $39.83_{\pm{0.21}}$
Peri-LN 3.279 $\textbf{2.614}_{\pm{0.011}}$ $3.178_{\pm{0.132}}$ $\textit{36.80}_{\pm{0.27}}$ $\textbf{67.05}_{\pm{0.12}}$ $\textbf{37.13}_{\pm{0.09}}$ $\textbf{31.44}_{\pm{0.61}}$ $49.06_{\pm{0.40}}$ $22.94_{\pm{0.02}}$ $52.75_{\pm{1.54}}$ $42.45_{\pm{0.28}}$
LNS 3.271 $2.652_{\pm{0.011}}$ $\textit{3.157}_{\pm{0.132}}$ $34.49_{\pm{0.17}}$ $65.59_{\pm{0.16}}$ $\textit{37.12}_{\pm{0.09}}$ $28.96_{\pm{0.54}}$ $50.61_{\pm{0.55}}$ $22.99_{\pm{0.02}}$ $52.69_{\pm{0.66}}$ $41.78_{\pm{0.20}}$
DyT 3.855 $3.361_{\pm{0.010}}$ $3.971_{\pm{0.132}}$ $29.78_{\pm{0.40}}$ $58.42_{\pm{0.53}}$ $25.95_{\pm{0.17}}$ $28.28_{\pm{0.90}}$ $49.53_{\pm{0.79}}$ $\textbf{23.46}_{\pm{0.13}}$ $\textit{56.64}_{\pm{1.02}}$ $38.87_{\pm{0.17}}$
BHyT 3.254 $2.693_{\pm{0.012}}$ $\textbf{3.130}_{\pm{0.133}}$ $34.61_{\pm{0.20}}$ $66.47_{\pm{0.27}}$ $36.95_{\pm{0.12}}$ $29.68_{\pm{0.46}}$ $\textit{51.14}_{\pm{1.02}}$ $\textit{23.07}_{\pm{0.05}}$ $\textbf{58.21}_{\pm{0.81}}$ $\textbf{42.88}_{\pm{0.19}}$
4.3 Instruction Fine-tuning and Benchmark Evaluation
We apply LoRA-based supervised fine-tuning to assess whether the stability achieved during pretraining carries over to instruction-tuned performance. Table 2 summarizes the SFT losses and downstream benchmark accuracies for the Llama-3B. Across both scales, BHyT maintains the stable optimization observed during pretraining and transfers this advantage effectively to the SFT stage. Notably, BHyT attains lower SFT loss and competitive accuracy on a broad set of benchmarks, demonstrating that its depth-wise stability and controlled activation dynamics continue to support reliable adaptation beyond pretraining. We present the pretraining and SFT results for the 1B model in Appendix E.